# Analysis and Synthesis Denoisers

### v2

@@ 5.2.2 Simulation results
Additionally, we give in Figure 7 reconstruction results for two images. The regularization parameters are chosen according to Figure 5 andFigure 6. For AD we chose $\lambda=10^{-3}$ that achieves best results for all testing configurations. For SD we chose $\lambda=10^{-2}$ that looks a compromise between stability and results (for all configurations). It is to be noted however that for $(L_{\text{train}},L_{\text{PnP}})=(1,1)$ and $(L_{\text{train}},L_{\text{PnP}})=(20,1)$, SD does not achieve convergence within the $10^{3}$ iterations, and results displayed in Figure 7 hence should be seen as early stopping results in this case. Finally, for DRUnet we take $\lambda=10^{-1}$ that is the only value for which DRUnet is stable [^2] Note that $\lambda=10^{-2}$ gives better PSNR values for DRUnet, however to the price of important artifact in the reconstruction.. Overall it can be observed that AD leads to the best reconstructions, with any configuration $(L_{\text{train}},L_{\text{den}})\in\{1,20\}^{2}$. Further, for fixed dictionary, doing $L_{\text{PnP}}=1$ or $L_{\text{PnP}}=20$ sub-iterations leads to the same reconstructions in our examples. This observation is not true for SD due to early stopping of the algorithms. In particular, SD produces artifacts at the edges of the images for both examples, even when high PSNR values are achieved.
$\overline{x}$
$y$
    $\lambda=10^{-5}$ $\lambda=10^{-3}$ $\lambda=10^{-2}$ $\lambda=10^{-1}$ SD $(L_{\text{train}},L_{\text{PnP}})=(1,20)$ Refer to caption Refer to caption Refer to caption Refer to caption PSNR=$17.35$dB PSNR=$22.54$dB PSNR=$23.60$dB PSNR=$26.12$dB AD $(L_{\text{train}},L_{\text{PnP}})=(1,20)$ Refer to caption Refer to caption Refer to caption Refer to caption PSNR=$15.69$dB PSNR=$27.42$dB PSNR=$22.14$dB PSNR=$8.15$dB DRUnet Refer to caption Refer to caption Refer to caption Refer to caption PSNR=$7.13$dB PSNR=$7.18$dB PSNR=$27.00$dB PSNR=$25.13$dB
@@
Figure 8: Example of FB-PnP outputs obtained with SD for $(L_{\text{train}},L_{\text{PnP}})=(1,20)$ (top row), AD for $(L_{\text{train}},L_{\text{PnP}})=(1,20)$ (middle row) and DRUnet (bottom row) when varying the regularization parameter $\lambda\in\{10^{-5},10^{-3},10^{-2},10^{-1}\}$.
@@
Finally, for the sake of completeness, we provide in Figure 8 reconstructed images obtained with the FB-PnP when varying $\lambda\in\{10^{-5},10^{-3},10^{-2},10^{-1}\}$. We show images obtained with SD and AD for $(L_{\text{train}},L_{\text{PnP}})=(1,20)$ (top and middle rows), and with DRUnet (bottom row). Interestingly it seems that SD is less sensitive to the choice of $\lambda$ than AD and DRUnet.
@@ 6 Conclusion
In this work we study convergence behavior of the FB-PnP algorithm where the proximity operator is replaced by a specific unfolded denoiser. In particular, we study two strategies: when the denoising problem is solved with an analysis formulation or when it is solved with a synthesis formulation. The two resulting analysis and the synthesis denoising minimization problems can then be solved with sub-iterations based on the dual FB algorithm and the FB algorithm, respectively. For both strategies, we show that computing only one sub-iteration within the FB-PnP combined with a warm-restart procedure on the sparse coefficients is equivalent to the FB-PnP algorithm when the full analysis or the synthesis denoising minimization problems are solved at each iteration. We further show that when smoothing the denoising problem using a Moreau envelope, this equivalence extends to an arbitrary number of sub-iterations. Finally, we investigated how these different strategies compare on a toy compressive sensing problem as well as on a deblurring imaging problem.

### v3

@@ 5.2.2 Simulation results
Additionally, we give in Figure 7 reconstruction results for two images. The regularization parameters are chosen according to Figure 5 and Figure 6. For AD we chose $\lambda=3\times 10^{-3}$ that achieves best results for all testing configurations. For SD we chose $\lambda=3\times 10^{-2}$ for $(L_{\text{train}},L_{\text{PnP}})\in\{(1,1),(20,1),(20,20\}$, and $\lambda_{\textsf{SD}}=10^{-2}$ for $(L_{\text{train}},L_{\text{PnP}})=(1,20)=10^{-10}$. Finally, for DRUnet we take $\lambda=10^{-1}$ that is the only value for which DRUnet is stable [^2] Note that $\lambda=3\times 10^{-2}$ gives better PSNR values for DRUnet, however to the price of important artifact in the reconstruction due to instability of the PnP iterations (see also Figure 8).. Overall it can be observed that AD leads to the best reconstructions, with any configuration $(L_{\text{train}},L_{\text{den}})\in\{1,20\}^{2}$. Further, for fixed dictionary, doing $L_{\text{PnP}}=1$ or $L_{\text{PnP}}=20$ sub-iterations leads to the same reconstructions in our examples. This observation is not true for SD due to early stopping of the algorithms. In particular, SD produces artifacts at the edges of the images for both examples, even when high PSNR values are achieved.
@@
Figure 8: Example of FB-PnP outputs obtained with SD for $(L_{\text{train}},L_{\text{PnP}})=(1,20)$ (top row), AD for $(L_{\text{train}},L_{\text{PnP}})=(1,20)$ (middle row) and DRUnet (bottom row) when varying the regularization parameter $\lambda\in\{10^{-3},3\times 10^{-3},3\times 10^{-2},10^{-1}\}$.
@@
Finally, for the sake of completeness, we provide in Figure 8 reconstructed images obtained with the FB-PnP when varying $\lambda\in\{10^{-3},3\times 10^{-3},3\times 10^{-2},10^{-1}\}$. We show images obtained with SD and AD for $(L_{\text{train}},L_{\text{PnP}})=(1,20)$ (top and middle rows), and with DRUnet (bottom row). Interestingly, with visual inspection, it seems that SD is less sensitive to the choice of $\lambda$ than AD and DRUnet, while AD is the more stable with respect to PSNR (despite very blurry reconstructions for large values of $\lambda$).
@@ 6 Conclusion
In this work we study convergence behavior of the FB-PnP algorithm where the proximity operator is replaced by a specific unfolded denoiser. In particular, we study two strategies: when the denoising problem is solved with an analysis formulation or when it is solved with a synthesis formulation. The two resulting analysis and the synthesis denoising minimization problems can then be solved with sub-iterations based on the dual FB algorithm and the FB algorithm, respectively. For both strategies, we show that computing only one sub-iteration within the FB-PnP combined with a warm-restart procedure on the sparse coefficients is equivalent to the FB-PnP algorithm when the full analysis or the synthesis denoising minimization problems are solved at each iteration. We further show that when smoothing the denoising problem using a Moreau envelope, under theoretical conditions, this equivalence extends to an arbitrary number of sub-iterations. Finally, we investigated how these different strategies compare on a toy compressive sensing problem as well as on a deblurring imaging problem in the context of DDL. Interestingly, we observed that training the AD with only $1$ iteration ($L_{\text{train}}=1$) or a larger number of iterations ($L_{\text{train}}=20$) does not seem to change much the denoising performances (see Figure 3). Further, for the FB-PnP, training with only $L_{\text{train}}=1$ iterations for both the SD and the AD leads to the best performances (see Figure 7). This suggests that choosing $L_{\text{train}}=1$ enables a fast and efficient denoiser training. In particular, AD with $L_{\text{train}}=1$ appeared to lead to the best results overall in our experiments.
Appendix A Loris-Verhoeven primal-dual algorithm
With our notation, the scaled Loris-Verhoeven primal-dual algorithm proposed in [38, Section 6, eq. (40)] is designed to $\text{find }\widehat{x}\in\underset{x\in\mathbb{R}^{N}}{\operatorname{Argmin}}\;\frac{1}{2}\|Ax-y\|^{2}+g_{\lambda}(\Gamma x)$, which is the same problem as stated in (2.6). The iterations to solve this problem are given by
[…]
with $\tau<2/\|A^{*}A\|$ and $\sigma<1/\|\Gamma^{*}\Gamma\|$.
Appendix B FB-PnP with DRUnet
DRUnet is a noise-aware denoising network that has been proposed in [61]. It has a U-Net like structure, with convolutional blocks in the encoder and decoder parts, and it takes into account the noise level of the input image as an extra input channel. It is considered as a state-of-the-art denoising network to use in PnP frameworks, and trained DRUnet can be found on PyTorch-based libraries such as DeepInv. Then, if $G\colon\mathbb{R}^{N}\times\mathbb{R}^{N}\to\mathbb{R}^{N}$ denotes the DRunet operator, the corresponding FB-PnP iterations are given as
[…]
where $\lambda>0$ denotes the input noise level that acts as a regularization parameter on the network, and $\tau>0$ is the step-size chosen to ensure convergence of $(x_{k})_{k\in\mathbb{N}}$.